Question 49
ProbabilityA box contains 5 blue balls, 3 red balls and 2 white balls. Two balls are selected from the box with replacement. Find the probability of obtaining two blue or two red balls.
- A.17/50
- B.3/25
- C.1/50
- D.3/250
Show the answer and explanation
Answer: A (17/50)
Explanation
P(two blue) = (5/10)2 = 1/4 and P(two red) = (3/10)2 = 9/100. Total = 25/100 + 9/100 = 34/100 = 17/50.